there are 132 packages to be delivered . one quarter of them will be delivered
to day. one third of the remaining will be delivered tomorrow. how many will be left to be delivered on the day after tomorrow ?
a) 44 b) 66 c) 22 d) 11 e) 33

Answer by alicealc(293)

today = 1/4 * 132 = 132/4 = 33
so,the remaining will be: 132 - 33 = 99

tomorrow = 1/3 * 99 = 99/3 = 33
so, the remaining will be: 99 - 33 = 66 (b)